# DQ-LoRe: Dual Queries with Low Rank Approximation Re-ranking for In-Context Learning

### v5

@@ 5 Conclusion
Appendix A Acknowledgement
This work is supported in part by National Key R&D Program of China under Grant No. 2020AAA0109700, Guangdong Outstanding Youth Fund (Grant No. 2021B1515020061), National Natural Science Foundation of China (NSFC)under Grant No.61976233, Mobility Grant Award under Grant No. M-0461, Shenzhen Science and Technology Program (Grant No. GJHZ20220913142600001), Nansha Key RD Program under Grant No.2022ZD014, National Natural Science Foundation of China under Grant No.62006255. We thank MindSpore [^1] https://www.mindspore.cn/ for the partial support of this work, which is a new deep learning computing framwork.
Appendix B Other Related Works
@@
Inspired by Li et al. (2020); Su et al. (2021); Huang et al. (2021), after obtaining a set of M exemplars, we perform dimensionality reduction on the embeddings of these M exemplars. We employ PCA Wold et al. (1987), as we believe it can effectively extract key reasoning information from the CoT to differentiate between different exemplars. Furthermore, Devlin et al. (2018) observes that reducing the 768-dimensional embeddings to 256-dimensional vectors from the BERT model could alleviate the issue of anisotropy in the vector space, where vectors are unevenly distributed and confined within a narrow conical space. This finding suggests that unsupervised training of sentence embeddings produces "universal" representations that contain many redundant features for specific domain applications. Removing these redundant features often leads to a dual benefit of enhanced speed and effectiveness (Wang et al., 2023b). Additionally, there are other methods for dimensionality reduction, such as Singular Value Decomposition (SVD) or low-rank approximation techniques (Ye, 2004; Hyvarinen, 1999; Pan et al., 2022). Furthermore, parameter-based training methods for dimensionality reduction have been proposed Hu et al. (2021). In contrast to the methods mentioned above, considering the advantages of not requiring training and having lower computational overhead, we opted for employing PCA for low-rank approximations.
@@
Appendix C Introduction of Our Baselines
@@
Appendix D Experiments on Commonsense Reasoning Datasets
In this section, we present experiments on commonsense reasoning datasets such as StrategyQA (Geva et al., 2021) and QASC (Khot et al., 2020) using GPT-3.5-Turbo-16k. The detailed experimental results are presented in the table below:
Table 6: Experiments on Question Answering Datasets.
Model StrategyQA QASC
CoT 73.8 81.8
Complex-CoT 74.5 75.8
Auto-CoT 71.2 74.1
EPR 73.4 80.2
CEIL 73.4 80.8
DQ-LoRe(PCA) 74.6 81.2
DQ-LoRe(Gaussian kernel) 75.4 82.7
@@ Appendix D Experiments on Commonsense Reasoning Datasets
Appendix E Experiments on More LLMs
In this section, we present experiments on a broader set of LLMs using the GSM8K dataset. The results are presented in the table below. Due to computational resource constraints, our experiments are primarily conducted on three 7 billion-parameter scale models: Llama2-7b-hf Touvron et al. (2023c), Llama2-7b-chat-hf Touvron et al. (2023c), and Vicuna-7b Chiang et al. (2023).
Table 7: Experiments on 7B-Scale Language Models on the GSM8K Dataset.
Model Llama2-7b-hf Llama2-7b-chat-hf Vicuna-7b
CoT 12.6 22.9 19.4
Complex-CoT 17.7 29.4 23.8
Auto-CoT 15.0 27.0 23.9
EPR 15.1 23.0 22.0
CEIL 15.2 26.7 22.4
DQ-LoRe 16.0 28.9 23.8
@@ Appendix E Experiments on More LLMs
Appendix F Ablation Study
Table 8: Experiments with Different Encoders and Ablation of LoRe Across Baselines on GSM8K.
Method BERT-base RoBERTa-base
EPR 77.3 78.0
EPR + LoRe 77.0 77.8
CEIL 79.4 77.7
CEIL + LoRe 78.7 77.4
DQ-LoRe w/o LoRe 78.9 76.4
DQ-LoRe 80.7 78.8
@@ Appendix F Ablation Study
To investigate the impact of different encoder models on retrieval performance, we examine the effects of employing various base models on the experimental results using GPT-3.5-Turbo-16k. Our findings indicate that, in terms of overall performance, the ’BERT-base’ model outperforms the ’RoBERTa-base’ (Liu et al., 2019) model, except for an enhancement observed in EPR when employing the ’RoBERTa-base’ model. Furthermore, we assess the influence of LoRe on various models and note that the adoption of LoRe leads to a decline in performance for both EPR and CEIL. This decline can primarily be attributed to the lack of the dual queries process, which impedes the incorporation of additional CoT information. As demonstrated in Section 4.4, conducting LoRe without directly encoding additional CoT information leads to diminished in-context learning performance.
Table 9: Ablation Study: Impact of Various Regularization Techniques in Re-Ranking Algorithm
Re-ranking method GSM8K SVAMP AQUA StrategyQA QASC
w/o dual queries 77.0 86.1 54.9 72.8 80.1
w/o re-ranking 78.9 89.0 57.0 74.1 82.1
Gaussian kernel 79.0 87.6 56.2 75.4 82.7
MGS 79.7 87.3 57.8 73.2 81.0
ICA 79.5 85.7 57.0 74.2 82.2
PCA 80.7 90.0 59.8 74.6 81.2
BERT-base 80.7 90.0 59.8 75.4 82.7
RoBERTa-base 78.8 89.0 57.8 74.2 81.2
@@
To present a comprehensive analysis of the impact of different regularization methods on the embedding space, we have executed additional experiments across various datasets. In the table 9, the ’w/o dual queries’ denotes the scenario where re-ranking is performed without engaging the query language model, a practice that has been empirically shown to detrimentally affect model performance. It highlights the significance of incorporating dual queries mechanisms to enhance the model’s ability to discern and prioritize CoT information effectively. The ’w/o re-ranking’ denotes scenario where no re-ranking process is implemented. It implies that the algorithm conducts a singular sorting operation following the query to LLMs. The term ’ICA’ (Independent Component Analysis) (Hyvärinen & Oja, 2000) represents a theoretically sound approach to achieving disentangled representations, which underscores the capacity to discern latent variables within the data. In this method, we reduce the feature dimensions of the embedding to 16 dimensions. In the table 9, ’BERT-base’ and ’RoBERTa-base’ are denoted as employing distinct encoders for initial-stage ranking similarity computations, and it’s noteworthy that they utilize the Gaussian kernel as the regularization method specifically for the StrategyQA and QASC datasets.
@@
Table 10: Comparing the Using of Trained Encoder and Solely CoT Similarity Retrieval.
Method w/o trainning with trainning
question + CoT 77.7 78.9
question + CoT with LoRe 78.8 80.7
only CoT 76.1 79.4
only CoT with LoRe 78.1 80.4
@@
In Table 10, we explore the impact of various encoder training strategies and ranking methodologies on the outcome. The "w/o training" configuration utilizes a BERT model, which has not been trained on a specific task, for encoding sample representations. Conversely, the "with training" setup involves a BERT model fine-tuned on a training dataset, aimed at enhancing encoding effectiveness through contrastive learning. The "question + CoT" scenario leverages both the query context and CoT insights simultaneously for encoding, while the "only CoT" approach focuses solely on encoding based on CoT information, establishing direct similarity measures between the CoT in the exemplars and the question’s CoT.
A comparative analysis between the "w/o training" and "with training" conditions reveals that encoder training on pertinent datasets markedly improves retrieval performance. Especially notable in the "only CoT" setting, training the encoder results in a 3.3% performance improvement. Our proposed method demonstrates resilience, consistently enhancing outcomes even without dataset-specific encoder training. Remarkably, in the "only CoT with LoRe" configuration under "w/o training," we attain a 2% absolute performance boost, underscoring the efficacy of our approach without necessitating specialized encoder training. Lastly, our findings suggest that relying solely on CoT similarity for encoding is less effective compared to a hybrid approach that incorporates both question and CoT, indicating the added value of integrating comprehensive contextual cues in the encoding process.
Table 11: Average Processing Time per Question for DQ-LoRe and Baseline.
Stage EPR CEIL DQ-LoRe
All Time 0.756s 0.758s 1.584s
First Query - - 0.728s
Second Query 0.029 0,032s 0.028s
Re-Ranking - - 0.101s
Inference 0.727s 0.726s 0.727s
@@
We conduct thorough tests to compare the time efficiency of our approach against baseline methods. Subsequently, we conducted a detailed analysis of the time consumption for each specific module within our method. For a more comprehensive understanding, please refer to the detailed experimental results provided in the subsequent table.
From Table 11, it is evident that the time expended by our method exhibits a linear relationship in comparison to the baseline. Further insights from Table 11 reveal that the predominant additional consumption, in contrast to the baseline, occurs during the initial stage of requesting the LLMs to acquire CoTs. In practice, each API key for the GPT-3.5-Turbo-16k model can process 180,000 tokens per minute, and with an average request requiring around 2000 tokens, theoretically, by employing numerous keys and parallelizing threads, the time consumption in this stage can be significantly reduced.
Appendix G Case Study on SVAMP
In this section, we present exemplars retrieved for a single data point on the SVAMP dataset. Tables 12, 13, 14, 15, and 16 respectively display the 8-shot results retrieved by EPR, CEIL, and DQ-LoRe. It can be observed that EPR relies on the pattern of word co-occurrence for retrieval, while CEIL retrieves some unrelated examples. From Tables 12, 13, it is visually evident that although EPR achieves the highest accuracy on SVAMP, it relies on patterns of word co-occurrence and happens to find the right prompt as a shortcut. This observation reveals that EPR has learned the phenomenon of spurious correlations. The model exhibits a tendency to group together all exemplars featuring word co-occurrence, which is prevalent in the SVAMP dataset due to its numerous analogous questions that mainly differ in numerical values. Consequently, EPR consistently retrieves exemplars characterized by significant word co-occurrence and analogous problem-solving methodologies.
@@ Appendix G Case Study on SVAMP
Table 12: EPR Case Study on SVAMP: Part 1
  Problem
@@
Question: Mary is baking a cake. The recipe calls for 6 cups of flour 8 cups of sugar and 7 cups of salt. She already put in 5 cups of flour.How many more cups of sugar than cups of salt does she need to add now?
EPR Prompt
Question: Mary is baking a cake. The recipe calls for 12 cups of flour 14 cups of sugar and 7 cups of salt. She already put in 2 cups of flour.How many more cups of flour than cups of salt does she need to add now?
A: Let’s think step by step To find out how many more cups of flour than cups of salt Mary needs to add now, we first need to calculate the total amount of flour and salt needed for the recipe. The recipe calls for 12 cups of flour and Mary has already added 2 cups, so she still needs 12 - 2 = 10 cups of flour. The recipe also calls for 7 cups of salt and Mary has not added any yet, so she needs 7 cups of salt. Therefore, Mary needs to add 10 - 7 = 3 more cups of flour than cups of salt. The answer is 3.
Question: Mary is baking a cake. The recipe calls for 9 cups of sugar 7 cups of flour and 4 cups of salt. She already put in 2 cups of flour.How many more cups of flour than cups of salt does she need to add now?
A: Let’s think step by step Mary needs to add 7 - 2 = 5 more cups of flour. She needs to add 5 - 4 = 1 more cup of flour than cups of salt. Answer: 111. The answer is 1.

### v6

@@ Acknowledgments
Appendix A Other Related Works
@@
Inspired by Li et al. (2020); Su et al. (2021); Huang et al. (2021), after obtaining a set of M exemplars, we perform dimensionality reduction on the embeddings of these M exemplars. We employ PCA Wold et al. (1987), as we believe it can effectively extract key reasoning information from the CoT to differentiate between different exemplars. Furthermore, Devlin et al. (2018) observes that reducing the 768-dimensional embeddings to 256-dimensional vectors from the BERT model could alleviate the issue of anisotropy in the vector space, where vectors are unevenly distributed and confined within a narrow conical space. This finding suggests that unsupervised training of sentence embeddings produces “universal” representations that contain many redundant features for specific domain applications. Removing these redundant features often leads to a dual benefit of enhanced speed and effectiveness (Wang et al., 2023b). Additionally, there are other methods for dimensionality reduction, such as Singular Value Decomposition (SVD) or low-rank approximation techniques (Ye, 2004; Hyvarinen, 1999; Pan et al., 2022). Furthermore, parameter-based training methods for dimensionality reduction have been proposed Hu et al. (2021). In contrast to the methods mentioned above, considering the advantages of not requiring training and having lower computational overhead, we opted for employing PCA for low-rank approximations.
@@
Appendix B Introduction of Our Baselines
@@
Appendix C Experiments on Commonsense Reasoning Datasets
In this section, we present experiments on commonsense reasoning datasets such as StrategyQA (Geva et al., 2021) and QASC (Khot et al., 2020) using GPT-3.5-Turbo-16k. The detailed experimental results are presented in Figure 5.
CoTComplex-CoTAuto-CoTEPRCEILDQ-LoRe(PCA)DQ-LoRe(Gaussian)707075758080858573.873.874.574.571.271.273.473.473.473.474.674.675.475.481.881.875.875.874.174.180.280.280.880.881.281.282.782.7Accuracy (%)StrategyQAQASC
Figure 5: Experiments on Commonsense Reasoning Datasets using GPT-3.5-Turbo-16k. DQ-LoRe with Gaussian kernel achieves the best performance on both StrategyQA (75.4%) and QASC (82.7%), outperforming all baseline methods.
@@ Appendix C Experiments on Commonsense Reasoning Datasets
Appendix D Experiments on More LLMs
In this section, we present experiments on a broader set of LLMs using the GSM8K dataset. The results are presented in Figure 6. Due to computational resource constraints, our experiments are primarily conducted on three 7 billion-parameter scale models: Llama2-7b-hf Touvron et al. (2023c), Llama2-7b-chat-hf Touvron et al. (2023c), and Vicuna-7b Chiang et al. (2023).
CoTComplex-CoTAuto-CoTEPRCEILDQ-LoRe1010151520202525303012.612.617.717.7151515.115.115.215.2161622.922.929.429.42727232326.726.728.928.919.419.423.823.823.923.9222222.422.423.823.8Accuracy (%)Llama2-7b-hfLlama2-7b-chat-hfVicuna-7b
Figure 6: Experiments on 7B-Scale Language Models on the GSM8K Dataset. Complex-CoT achieves the best performance on Llama2-7b-hf (17.7%) and Llama2-7b-chat-hf (29.4%), while Auto-CoT performs best on Vicuna-7b (23.9%). DQ-LoRe shows competitive performance across all models, achieving 16.0%, 28.9%, and 23.8% respectively.
@@ Appendix D Experiments on More LLMs
Appendix E Ablation Study
EPREPR+LoReCEILCEIL+LoReDQ-LoRe w/o LoReDQ-LoRe7474767678788080828277.377.3777779.479.478.778.778.978.980.780.7787877.877.877.777.777.477.476.476.478.878.8Accuracy (%)BERT-baseRoBERTa-base
Figure 7: Experiments with different encoders and ablation of LoRe across baselines on GSM8K. Key findings: (1) Adding LoRe to EPR/CEIL slightly decreases performance, indicating LoRe alone is not beneficial without dual queries. (2) DQ-LoRe achieves the best results on both encoders (80.7% with BERT-base, 78.8% with RoBERTa-base). (3) The significant gap between “DQ-LoRe w/o LoRe” and “DQ-LoRe” demonstrates the effectiveness of LoRe when combined with dual queries.
@@ Appendix E Ablation Study
To investigate the impact of different encoder models on retrieval performance, we examine the effects of employing various base models on the experimental results using GPT-3.5-Turbo-16k. As shown in Figure 7, our findings indicate that, in terms of overall performance, the “BERT-base” model outperforms the “RoBERTa-base” (Liu et al., 2019) model, except for an enhancement observed in EPR when employing the “RoBERTa-base” model. Furthermore, we assess the influence of LoRe on various models and note that the adoption of LoRe leads to a decline in performance for both EPR and CEIL. This decline can primarily be attributed to the lack of the dual queries process, which impedes the incorporation of additional CoT information. As demonstrated in Section 4.4, conducting LoRe without directly encoding additional CoT information leads to diminished in-context learning performance.
Table 4: Ablation Study: Impact of Various Regularization Techniques in Re-Ranking Algorithm. Results are accuracy (%) using GPT-3.5-Turbo-16k. Best results for each dataset are in bold.
Category Method GSM8K SVAMP AQUA StrategyQA QASC Avg. $\Delta$ Ablation w/o dual queries 77.0 86.1 54.9 72.8 80.1 74.2 - w/o re-ranking 78.9 89.0 57.0 74.1 82.1 76.2 +2.0 Regularization Method Gaussian kernel 79.0 87.6 56.2 75.4 82.7 76.2 +2.0 MGS 79.7 87.3 57.8 73.2 81.0 75.8 +1.6 ICA 79.5 85.7 57.0 74.2 82.2 75.7 +1.5 PCA (Ours) 80.7 90.0 59.8 74.6 81.2 77.3 +3.1 Encoder BERT-base 80.7 90.0 59.8 75.4 82.7 77.7 +3.5 RoBERTa-base 78.8 89.0 57.8 74.2 81.2 76.2 +2.0
@@
To present a comprehensive analysis of the impact of different regularization methods on the embedding space, we have executed additional experiments across various datasets. In the table 4, the “w/o dual queries” denotes the scenario where re-ranking is performed without engaging the query language model, a practice that has been empirically shown to detrimentally affect model performance. It highlights the significance of incorporating dual queries mechanisms to enhance the model’s ability to discern and prioritize CoT information effectively. The “w/o re-ranking” denotes scenario where no re-ranking process is implemented. It implies that the algorithm conducts a singular sorting operation following the query to LLMs. The term “ICA” (Independent Component Analysis) (Hyvärinen and Oja, 2000) represents a theoretically sound approach to achieving disentangled representations, which underscores the capacity to discern latent variables within the data. In this method, we reduce the feature dimensions of the embedding to 16 dimensions. In the table 4, “BERT-base” and “RoBERTa-base” are denoted as employing distinct encoders for initial-stage ranking similarity computations, and it’s noteworthy that they utilize the Gaussian kernel as the regularization method specifically for the StrategyQA and QASC datasets.
@@
question+CoTquestion+CoT+LoReonly CoTonly CoT+LoRe7474767678788080828277.777.778.878.876.176.178.178.178.978.980.780.779.479.480.480.4Accuracy (%)w/o trainingwith training
Figure 8: Comparing the impact of trained encoder and CoT similarity retrieval on GSM8K using GPT-3.5-Turbo-16k. Key findings: (1) Training the encoder consistently improves performance across all methods (+1.2% to +3.3%). (2) Adding LoRe improves both “question+CoT” (+1.1% w/o training, +1.8% with training) and “only CoT” (+2.0% w/o training, +1.0% with training). (3) The best result (80.7%) is achieved by combining question+CoT with LoRe and trained encoder.
@@
In Figure 8, we explore the impact of various encoder training strategies and ranking methodologies on the outcome. The “w/o training” configuration utilizes a BERT model, which has not been trained on a specific task, for encoding sample representations. Conversely, the “with training” setup involves a BERT model fine-tuned on a training dataset, aimed at enhancing encoding effectiveness through contrastive learning. The “question + CoT” scenario leverages both the query context and CoT insights simultaneously for encoding, while the “only CoT” approach focuses solely on encoding based on CoT information, establishing direct similarity measures between the CoT in the exemplars and the question’s CoT.
A comparative analysis between the “w/o training” and “with training” conditions reveals that encoder training on pertinent datasets markedly improves retrieval performance. Especially notable in the “only CoT” setting, training the encoder results in a 3.3% performance improvement. Our proposed method demonstrates resilience, consistently enhancing outcomes even without dataset-specific encoder training. Remarkably, in the “only CoT with LoRe” configuration under “w/o training”, we attain a 2% absolute performance boost, underscoring the efficacy of our approach without necessitating specialized encoder training. Lastly, our findings suggest that relying solely on CoT similarity for encoding is less effective compared to a hybrid approach that incorporates both question and CoT, indicating the added value of integrating comprehensive contextual cues in the encoding process.
[…]
Second QueryRe-RankingInferenceTotal00.50.5111.51.5Time (seconds)EPRCEILDQ-LoRe
Figure 9: Average processing time per question for DQ-LoRe and baselines. Left: Stacked bar chart showing time breakdown by stage. Right: Line chart comparing methods across stages. DQ-LoRe requires additional time for First Query (0.728s) and Re-Ranking (0.101s), resulting in 1.584s total vs $\sim$0.76s for baselines. However, the inference time remains identical (0.727s), and the extra time is justified by the significant accuracy improvements.
@@
We conduct thorough tests to compare the time efficiency of our approach against baseline methods. Subsequently, we conducted a detailed analysis of the time consumption for each specific module within our method. For a more comprehensive understanding, please refer to the detailed experimental results provided in Figure 9.
From Figure 9, it is evident that the time expended by our method exhibits a linear relationship in comparison to the baseline. Further insights from Figure 9 reveal that the predominant additional consumption, in contrast to the baseline, occurs during the initial stage of requesting the LLMs to acquire CoTs. In practice, each API key for the GPT-3.5-Turbo-16k model can process 180,000 tokens per minute, and with an average request requiring around 2000 tokens, theoretically, by employing numerous keys and parallelizing threads, the time consumption in this stage can be significantly reduced.
Appendix F Case Study on SVAMP
In this section, we present exemplars retrieved for a single data point on the SVAMP dataset. Tables F, F, F, F, and F respectively display the 8-shot results retrieved by EPR, CEIL, and DQ-LoRe. It can be observed that EPR relies on the pattern of word co-occurrence for retrieval, while CEIL retrieves some unrelated examples. From Tables F, F, it is visually evident that although EPR achieves the highest accuracy on SVAMP, it relies on patterns of word co-occurrence and happens to find the right prompt as a shortcut. This observation reveals that EPR has learned the phenomenon of spurious correlations. The model exhibits a tendency to group together all exemplars featuring word co-occurrence, which is prevalent in the SVAMP dataset due to its numerous analogous questions that mainly differ in numerical values. Consequently, EPR consistently retrieves exemplars characterized by significant word co-occurrence and analogous problem-solving methodologies.
@@ Appendix F Case Study on SVAMP
Test Problem Question: Mary is baking a cake. The recipe calls for 6 cups of flour, 8 cups of sugar and 7 cups of salt. She already put in 5 cups of flour. How many more cups of sugar than cups of salt does she need to add now? Answer: 8 - 7 = 1 more cup of sugar than salt.
EPR Retrieved Exemplars Observation: EPR retrieves exemplars with high word co-occurrence (all about “Mary baking a cake”), exploiting spurious correlations in SVAMP. Exemplar 1: Mary is baking a cake. The recipe calls for 12 cups of flour, 14 cups of sugar and 7 cups of salt. She already put in 2 cups of flour. How many more cups of flour than cups of salt does she need to add now? A: The recipe calls for 12 cups of flour and Mary has already added 2 cups, so she still needs 12 - 2 = 10 cups of flour. The recipe also calls for 7 cups of salt. Therefore, Mary needs to add 10 - 7 = 3 more cups of flour than cups of salt. Exemplar 2: Mary is baking a cake. The recipe calls for 9 cups of sugar, 7 cups of flour and 4 cups of salt. She already put in 2 cups of flour. How many more cups of flour than cups of salt does she need to add now? A: Mary needs to add 7 - 2 = 5 more cups of flour. She needs to add 5 - 4 = 1 more cup of flour than cups of salt. Exemplar 3: Mary is baking a cake. The recipe calls for 14 cups of flour and 6 cups of sugar. She already put in 5 cups of flour. How many more cups of flour than cups of sugar does she need to add now? A: Mary needs to add 14 - 5 = 9 more cups of flour. She needs to add 9 - 6 = 3 more cups of flour than cups of sugar. Exemplar 4: Mary is baking a cake. The recipe calls for 11 cups of flour and 7 cups of sugar. She already put in some cups of flour. If she still needs 2 more cups of flour than sugar, how many cups of flour did she put in? A: Let x = cups of flour already put in. She needs (7 + 2) = 9 more cups of flour. So x + 9 = 11, thus x = 2. Exemplar 5: Mary is baking a cake. The recipe calls for 3 cups of sugar, 10 cups of flour and 15 cups of salt. She already put in 6 cups of flour. How many more cups of flour does she need to add? A: Total flour required is 10 cups, Mary already put in 6 cups. She needs to add 10 - 6 = 4 more cups of flour.